OpenStudy (anonymous):

Find an equation of the tangent line to the curve at the point (4,2)

OpenStudy (anonymous):

\[\LARGE f(x)=\frac{x-2}{x-3}\]

OpenStudy (amistre64):

the slope of a tangent line if f' use the point slope format to solve the for the line: y - y1 = m(x-x1) such that m = f'(4) and x1,y1 is the given point